# CGKN: A Deep Learning Framework for Modeling Complex Dynamical Systems and Efficient Data Assimilation

### v1

@@ 1 Introduction
Data assimilation (DA) [46, 47, 48, 49, 50, 51] integrates observational data into models to enhance state estimation. DA was originally designed to improve the estimation of initial conditions for the real-time forecast. Nowadays, DA has wide applications in parameter estimation, model identification, optimal control, and reconstructing missing data. DA consists of a two-step procedure: a model forecast generates a statistical prediction, which is then refined in the analysis step by incorporating new observations, improving accuracy, and reducing uncertainty. Given the highly nonlinear nature of many complex dynamical systems, ensemble DA has become one of the most widely used methods in practice [52, 53, 54, 55]. When the governing equations of a dynamical system are unknown or too costly to simulate, machine learning can often assist DA in two ways. First, machine learning can be used to build data-driven forecast models for DA [56, 57, 58, 59, 60, 61, 62, 63]. Second, machine learning can streamline the entire DA process by developing end-to-end learning frameworks [64, 65, 66, 67, 68, 69]. A comprehensive review of machine learning with data assimilation can be found in [70].
@@
In this paper, we develop a Conditional Gaussian Koopman Network (CGKN), which is a deep learning framework designed to address the challenges of modeling complex dynamical systems, performing efficient forecasts, and accelerating effective DA. The CGKN is structured as a neural stochastic differential equation. It operates in three main steps: (i) transforming the state of the original nonlinear system into that of a conditional Gaussian nonlinear system, (ii) learning the dynamics of the conditional Gaussian nonlinear system to enable rapid state forecasting and DA, and (iii) transforming the forecast and DA results back into the original system. Notably, the transformations required in the first and third steps are often unknown a priori, but they can be jointly learned along with the dynamics of the conditional Gaussian nonlinear system.
@@ 2 Methodology
where $\mathbf{u}\in\mathbb{R}^{d_{\mathbf{u}}}$ denotes the system state and $\mathcal{M}:\mathbb{R}^{d_{\mathbf{u}}}\mapsto\mathbb{R}^{d_{\mathbf{u}}}$ is a nonlinear function that describes the system dynamics. In this work, we focus on partially observable dynamical systems, with observed state variables denoted as $\mathbf{u}_{1}\in\mathbb{R}^{d_{\mathbf{u}_{1}}}$ and the unobserved state variables denoted as $\mathbf{u}_{2}\in\mathbb{R}^{d_{\mathbf{u}_{2}}}$. To explicitly discriminate the observed state variables $\mathbf{u}_{1}$ and unobserved state variables $\mathbf{u}_{2}$, the true system in Eq. (2.1) is rewritten as:
@@
where $\mathcal{M}_{1}:\mathbb{R}^{d_{\mathbf{u}}}\mapsto\mathbb{R}^{d_{\mathbf{u}_{1}}}$ and $\mathcal{M}_{2}:\mathbb{R}^{d_{\mathbf{u}}}\mapsto\mathbb{R}^{d_{\mathbf{u}_{2}}}$. The DA task targets at the conditional distribution $p(\mathbf{u}_{2}(t)|\mathbf{u}_{1}(s),s\leq t)$, where the trajectory of $\mathbf{u}_{1}$ up to time $t$, namely $\mathbf{u}_{1}(s\leq t)$, is the observations while the state of $\mathbf{u}_{2}$ at $t$ needs to be estimated [46, 49, 47].
@@
Both $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ can be strongly nonlinear functions, which can be approximated by neural networks according to universal approximation theory [83, 84, 85]. With a sufficient amount of data, the approximation can be quite accurate, leading to a good performance of state forecast. However, the DA would require techniques such as the ensemble-based method which are computationally expensive, could suffer from sampling errors, and often require empirical tuning.
@@
In this work, we develop the Conditional Gaussian Koopman Network (CGKN) for both state forecast and efficient DA. Instead of directly building a data-driven model for state variables $\mathbf{u}_{1}$ and $\mathbf{u}_{2}$, the proposed framework leverages a generalized version of the Koopman theory and exploits a proper transformation to map the unobserved state variables $\mathbf{u}_{2}$ to latent state variables $\mathbf{v}$, such that those latent variables remain nonlinear but are conditionally linear in the system dynamics. Therefore, the observed state variables $\mathbf{u}_{1}$ and the latent state variables $\mathbf{v}$ can be modeled by a conditional Gaussian nonlinear system [72, 73, 71], which encompasses many nonlinear models across various disciplines with many applications in natural science and engineering [86, 87, 88, 89, 90, 91, 92]. Importantly, the conditional Gaussian nonlinear system facilitates the development of closed analytic formulae for DA. The analytic formulae provide more efficient DA than those techniques developed for nonlinear models and also enable the incorporation of DA performance into the loss function when training the CGKN model. A schematic overview of the CGKN framework, including its application in state forecasting and DA, is presented in Fig. 2.1.
@@
Figure 2.1: Schematic diagram of conditional Gaussian Koopman network (CGKN) for a partially observable system. The complex nonlinear dynamical system is transformed into the conditional Gaussian nonlinear system via a generalized usage of Koopman theory, which maps between the unobserved state variables and the latent state variables featured by conditional linear dynamics. The unknown dynamics of the conditional Gaussian nonlinear system are constructed by neural networks and jointly learned with the unknown nonlinear mappings. The analytic DA formulae for the conditional Gaussian nonlinear system can significantly accelerate DA process and also allow a computationally affordable DA loss to be incorporated into the CGKN training process. The trained CGKN model can be used for state forecasts and efficient DA, for which the computations are mainly performed in the conditional Gaussian state space and then mapped back to the original state space.
@@ 2.1 Conditional Gaussian Koopman Network (CGKN)
In this work, the proposed CGKN approximates the true nonlinear system described in Eq. (LABEL:eq:DS_u1u2) and comprises three components: (i) an encoder that maps the original unobservable system state variables to the latent state variables of a conditional Gaussian nonlinear system, (ii) neural networks that approximate the unknown terms in the dynamics of the conditional Gaussian nonlinear system, and (iii) a decoder that transforms the latent state variables of the conditional Gaussian nonlinear system back to the original unobservable system state variables. The conditional Gaussian nonlinear system is written as:
@@
where $\mathbf{f}_{1}$, $\mathbf{g}_{1}$, $\mathbf{f}_{2}$, and $\mathbf{g}_{2}$ are all arbitrarily nonlinear functions of $\mathbf{u}_{1}$ while $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ are independent Wiener processes. The latent state variables are given by $\mathbf{v}=\bm{\varphi}(\mathbf{u}_{2})$ with a nonlinear encoder $\bm{\varphi}:\ \mathbb{R}^{d_{\mathbf{u}_{2}}}\mapsto\mathbb{R}^{d_{\mathbf{v}}}$. It is worth noting that the conditional distribution of the latent variables $p(\mathbf{v}(t)|\mathbf{u}_{1}(s),s\leq t)$ is Gaussian, due to the conditional linear structure in Eq. (2.3), which facilitates efficient analytic formulae of DA with more details discussed in Section 2.1.2. In practice, the nonlinear functions $\mathbf{f}_{1}$, $\mathbf{g}_{1}$, $\mathbf{f}_{2}$, $\mathbf{g}_{2}$ in Eq. (2.3) are often unknown and can be approximated by neural networks. With the state prediction and DA results of $(\mathbf{u}_{1},\mathbf{v})^{\mathtt{T}}$, the corresponding original unobserved state variables can be obtained via $\mathbf{u}_{2}=\bm{\psi}(\mathbf{v})$ with the nonlinear decoder $\bm{\psi}:\ \mathbb{R}^{d_{\mathbf{v}}}\mapsto\mathbb{R}^{d_{\mathbf{u}_{2}}}$. Considering that the proper encoder and decoder are often unknown a priori, they can be jointly learned with the unknown nonlinear functions $\mathbf{f}_{1}$, $\mathbf{g}_{1}$, $\mathbf{f}_{2}$, $\mathbf{g}_{2}$.
@@
Figure 2.2: The architecture of CGKN and its applications for state prediction and DA. The CGKN approximates the true nonlinear complex dynamical system by a modeled system with a conditional linear structure, which is also known as a conditional Gaussian nonlinear system. The conditional linear structure is enabled by a proper choice of latent state variables $\mathbf{v}$ as illustrated in (a). The nonlinear mappings between the original state variables $\mathbf{u}_{2}$ and the latent state variables $\mathbf{v}$ are achieved via the encoder $\bm{\varphi}$ and decoder $\bm{\psi}$, which are jointly learned with the unknown functions $\mathbf{f}_{1}$, $\mathbf{g}_{1}$, $\mathbf{f}_{2}$, $\mathbf{g}_{2}$ of the conditional Gaussian nonlinear system. The learned CGKN model can be used for state prediction and efficient DA as illustrated in (b) and (c).
@@ 2.1.1 CGKN for State Prediction
The CGKN model can be used for state prediction given an initial state $\mathbf{u}^{\dagger}(t_{0})=(\mathbf{u}_{1}^{\dagger}(t_{0}),\mathbf{u}^{\dagger}_{2}(t_{0}))^{\mathtt{T}}$. Firstly, the unobserved state variables $\mathbf{u}^{\dagger}_{2}(t_{0})$ are transformed into latent state variables $\mathbf{v}^{\dagger}(t_{0})$ via an encoder $\bm{\varphi}$, which leads to the initial state of conditional Gaussian nonlinear system $\mathbf{u}_{\textrm{CG}}=(\mathbf{u}_{1},\mathbf{v})^{\mathtt{T}}$ for Eq. (2.3):
$\mathbf{u}_{\textrm{CG}}^{\dagger}(t_{0})=\begin{bmatrix}\mathbf{u}^{\dagger}_{1}(t_{0})\\ \mathbf{v}^{\dagger}(t_{0})\end{bmatrix}=\begin{bmatrix}\mathbf{u}^{\dagger}_{1}(t_{0})\\ \bm{\varphi}(\mathbf{u}^{\dagger}_{2}(t_{0}))\end{bmatrix}.$ (2.4)
@@
$\mathbf{u}_{\textrm{CG}}(t_{n})=\begin{bmatrix}\mathbf{u}_{1}(t_{n})\\ \mathbf{v}(t_{n})\end{bmatrix}=\begin{bmatrix}\mathbf{u}^{\dagger}_{1}(t_{0})+\int_{t_{0}}^{t_{n}}(\mathbf{f}_{1}(\mathbf{u}_{1}(t))+\mathbf{g}_{1}(\mathbf{u}_{1}(t))\mathbf{v})\mathrm{d}t\\ \mathbf{v}^{\dagger}(t_{0})+\int_{t_{0}}^{t_{n}}(\mathbf{f}_{2}(\mathbf{u}_{1}(t))+\mathbf{g}_{2}(\mathbf{u}_{1}(t))\mathbf{v})\mathrm{d}t\end{bmatrix}.$ (2.5)
@@
In this work, the encoder $\bm{\varphi}$, the decoder $\bm{\psi}$, and the functions $\mathbf{f}_{1}$, $\mathbf{g}_{1}$, $\mathbf{f}_{2}$, $\mathbf{g}_{2}$ are jointly learned to ensure a good state prediction performance, i.e., the difference between the state prediction $\mathbf{u}(t_{n})$ and the true system state $\mathbf{u}^{\dagger}(t_{n})$ are small for a series of $t_{n}$. This is achieved by incorporating the state prediction performance into the loss function, and more details of the loss function for training a CGKN model are summarized in Section 2.2. An illustration of CGKN for state prediction is shown in Fig. 2.2(b).

### v2

@@ 1 Introduction
Data assimilation (DA) [46, 47, 48, 49, 50, 51] integrates observational data into models to enhance state estimation. DA was originally designed to improve the estimation of initial conditions for the real-time forecast. Nowadays, DA has wide applications in parameter estimation, model identification, optimal control, and reconstructing missing data. DA consists of a two-step procedure: a model forecast generates a statistical prediction, which is then refined in the analysis step by incorporating new observations, improving accuracy, and reducing uncertainty. Given the highly nonlinear nature of many complex dynamical systems, ensemble DA [52, 53, 54, 55] has become one of the most widely used methods in practice. Ensemble DA methods utilize an ensemble of realizations to approximate probabilistic distributions of target states and update the ensemble accordingly when new observations become available. When the governing equations of dynamical systems are unknown or too costly to simulate, machine learning is often employed to assist DA. First, machine learning can be used to build data-driven forecast models for DA [56, 57, 58, 59, 60, 61, 62, 63]. Second, machine learning with observers enhances the prediction of complex systems, and the physics-informed machine learning can be used to assist in the design of observers [64, 65, 66]. Third, machine learning can streamline the entire DA process by developing end-to-end learning frameworks [67, 68, 69, 70, 71, 72]. A comprehensive review of machine learning with data assimilation can be found in [73].
@@
In this paper, we develop a Conditional Gaussian Koopman Network (CGKN), which is a deep learning framework designed to address the challenges of modeling complex dynamical systems, performing efficient forecasts, and accelerating effective DA. The CGKN is structured as a neural stochastic differential equation. It operates in three main steps: (i) transforming the state of the original nonlinear system into that of a conditional Gaussian nonlinear system, (ii) learning the dynamics of the conditional Gaussian nonlinear system to enable rapid state forecast and DA, and (iii) transforming the forecast and DA results back into the original system. Notably, the transformations required in the first and third steps are often unknown a priori, but they can be jointly learned along with the dynamics of the conditional Gaussian nonlinear system.
@@ 2 Methodology
where $\mathbf{u}\in\mathbb{R}^{d_{\mathbf{u}}}$ denotes the system state and $\mathcal{M}:\mathbb{R}^{d_{\mathbf{u}}}\mapsto\mathbb{R}^{d_{\mathbf{u}}}$ is a real vector-valued analytic function that describes the system dynamics. The system state $\mathbf{u}$ can be a high-dimensional vector and the dynamics function $\mathcal{M}$ can be highly nonlinear. The system in the form of Eq. (2.1) can encompass a wide range of complex dynamical behaviors including multi-scale dynamics, chaos, turbulence, stochasticity, non-Gaussian statistics, intermittency, and extreme events. In this work, we focus on partially observed dynamical systems, with observed state variables denoted as $\mathbf{u}_{1}\in\mathbb{R}^{d_{\mathbf{u}_{1}}}$ and the unobserved state variables denoted as $\mathbf{u}_{2}\in\mathbb{R}^{d_{\mathbf{u}_{2}}}$. The unobserved states are the states that are not directly measured and to be inferred from data of observed states through techniques of data assimilation (DA). To explicitly discriminate the observed state variables $\mathbf{u}_{1}$ and unobserved state variables $\mathbf{u}_{2}$, the true system in Eq. (2.1) is rewritten as:
@@
where $[\mathbf{u}_{1};\mathbf{u}_{2}]=\mathbf{u}$, $\mathcal{M}_{1}:\mathbb{R}^{d_{\mathbf{u}}}\mapsto\mathbb{R}^{d_{\mathbf{u}_{1}}}$ and $\mathcal{M}_{2}:\mathbb{R}^{d_{\mathbf{u}}}\mapsto\mathbb{R}^{d_{\mathbf{u}_{2}}}$ are real vector-valued analytic functions. The notation [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] is used to denote the vertical concatenation of two column vectors. The DA task targets at the conditional distribution $p(\mathbf{u}_{2}(t)|\mathbf{u}_{1}(s),s\leq t)$, where the trajectory of $\mathbf{u}_{1}$ up to time $t$, namely $\mathbf{u}_{1}(s\leq t)$, is the observations while the state of $\mathbf{u}_{2}$ at $t$ needs to be estimated [46, 49, 47].
@@
As a fundamental component of deep learning, a neural network is a computational model inspired by the structure of biological neural networks, consisting of hierarchical layers of neurons, each computing a weighted sum of inputs followed by a non-linear activation function. The universal approximation theorem [86, 87, 88] states that a neural network with at least one hidden layer, a sufficient number of neurons, and a non-linear activation function can approximate any continuous function. Therefore, both $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$, which can be highly nonlinear functions, can be approximated by neural networks according to universal approximation theory. With a sufficient amount of data, the approximation of neural networks can be quite accurate, leading to a good performance of state forecasts. However, the DA for the black-box neural-networks-based model would require techniques such as the ensemble-based method, which are computationally expensive, could suffer from sampling errors, and often require empirical tuning.
@@
In this work, we develop the Conditional Gaussian Koopman Network (CGKN) for both state forecast and efficient DA. Instead of directly building a data-driven model for state variables $\mathbf{u}_{1}$ and $\mathbf{u}_{2}$, the proposed framework leverages a generalized version of the Koopman theory and exploits a proper transformation to map the unobserved state variables $\mathbf{u}_{2}$ to latent state variables $\mathbf{v}$, such that those latent variables are conditionally linear in the system dynamics. Therefore, the observed state variables $\mathbf{u}_{1}$ and the latent state variables $\mathbf{v}$ can be modeled by a conditional Gaussian nonlinear system [75, 76, 74], which encompasses many nonlinear models across various disciplines with many applications in natural science and engineering [89, 90, 91, 92, 93, 94, 95]. Importantly, the conditional Gaussian nonlinear system facilitates the development of closed analytic formulae for DA. The analytic formulae provide more efficient DA than those techniques developed for nonlinear models and also enable the incorporation of DA performance into the loss function when training the CGKN model. It is essential to highlight that the governing equations of complex dynamical systems do not need to be known as CGKN is a data-driven method based on deep learning. If the governing equations are available, CGKN can be learned from the synthetic data generated from the simulation of the complex system and then provides efficient data assimilation. Alternatively, even if the governing equations are unknown, CGKN can still be learned using real data, for which this work assumes the data of unobserved states is accessible at the offline training stage. With a pre-trained CGKN, the inference processes, including both state prediction and efficient DA, do not require data from unobserved states. A schematic overview of the CGKN framework, including its application in state forecast and DA, is presented in Fig. 2.1.
@@
Figure 2.1: Schematic diagram of conditional Gaussian Koopman network (CGKN) for a partially observed system. The complex nonlinear dynamical system is transformed into the conditional Gaussian nonlinear system via a generalized usage of Koopman theory, which maps between the unobserved state variables and the latent state variables featured by conditional linear dynamics. The unknown dynamics of the conditional Gaussian nonlinear system are constructed by neural networks and jointly learned with the unknown nonlinear mappings. The analytic DA formulae for the conditional Gaussian nonlinear system can significantly accelerate DA process and also allow a computationally affordable DA loss to be incorporated into the CGKN training process. The trained CGKN model can be used for state forecasts and efficient DA, for which the computations are mainly performed in the conditional Gaussian state space and then mapped back to the original state space.
2.1 Koopman Theory
Koopman theory [77] provides a linear perspective on nonlinear dynamical systems in the form of Eq. (2.1) by describing the system in terms of the evolution of observable of the system state. The observable $h:\mathbb{R}^{d{{}_{\mathbf{u}}}}\mapsto\mathbb{R}$ is a function in the Hilbert space operating on the system state $\mathbf{u}$. For the complex dynamical system in Eq. (2.1), the state transition operator $\mathcal{G}^{\Delta t}:\mathbb{R}^{d{{}_{\mathbf{u}}}}\mapsto\mathbb{R}^{{d_{\mathbf{u}}}}$ over time interval $\Delta t$ is defined as:
[…]
The Koopman operator $\mathcal{K}^{\Delta t}$ is a linear operator acting on observable function $h$ which is defined as
$\mathcal{K}^{\Delta t}(h):=h\circ\mathcal{G}^{\Delta t},$ (2.4)
where $\circ$ denotes function composition.
Furthermore, the Koopman infinitesimal generator $\mathcal{L}$ is defined as:
$\mathcal{L}:=\lim_{\Delta t\rightarrow 0}\frac{\mathcal{K}^{\Delta t}(h)-h}{\Delta t},$ (2.5)
which characterizes the $\mathcal{L}$ as the rate of change of the observable $h$ since $\mathcal{K}^{\Delta t}(h(\mathbf{u}(t)))=h\circ\mathcal{G}^{\Delta t}(\mathbf{u}(t))=h(\mathbf{u}(t+\Delta t))$.
Therefore, the evolution of observable $h$ is governed by the Koopman infinitesimal generator $\mathcal{L}$ which is a linear operator:
$\frac{\partial h}{\partial t}=\mathcal{L}h.$ (2.6)
In real application, the infinite-dimensional linear dynamical system and the linear operator $\mathcal{L}$ are approximated by the finite-dimensional representations:
$\frac{\mathrm{d}\mathbf{v}}{\mathrm{d}t}=\mathbf{A}\mathbf{v},$ (2.7)
where $\mathbf{v}\in\mathbb{R}^{d_{\mathbf{v}}}$ is a finite-dimensional vector that approximates the infinite-dimensional function $h$ and the matrix $\mathbf{A}\in\mathbb{R}^{d_{\mathbf{v}}\times d_{\mathbf{v}}}$ is the linear dynamics of this discrete linear dynamical system which approximates the linear operator $\mathcal{L}$.
@@ 2.2 Conditional Gaussian Koopman Network (CGKN)
In this work, a generalized application of Koopman theory has been employed to linearize the unobserved states $\mathbf{u}_{2}$ while retaining the nonlinearity of observed states $\mathbf{u}_{1}$ for the partially observed dynamical systems in Eq. (LABEL:eq:DS_u1u2). The resulting conditional linear dynamical system is called the Conditional Gaussian Nonlinear System (CGNS), which is written as:
@@
where $\mathbf{f}_{1}:\mathbb{R}^{d_{\mathbf{u}_{1}}}\mapsto\mathbb{R}^{d_{\mathbf{u}_{1}}}$, $\mathbf{g}_{1}:\mathbb{R}^{d_{\mathbf{u}_{1}}}\mapsto\mathbb{R}^{d_{\mathbf{u}_{1}}\times d_{\mathbf{v}}}$, $\mathbf{f}_{2}:\mathbb{R}^{d_{\mathbf{u}_{1}}}\mapsto\mathbb{R}^{d_{\mathbf{v}}}$, and $\mathbf{g}_{2}:\mathbb{R}^{d_{\mathbf{u}_{1}}}\mapsto\mathbb{R}^{d_{\mathbf{v}}\times d_{\mathbf{v}}}$ are nonlinear functions of $\mathbf{u}_{1}$ to be learned from data. $\mathbf{W}_{1}\in\mathbb{R}^{d_{\mathbf{u}_{1}}}$ and $\mathbf{W}_{2}\in\mathbb{R}^{d_{\mathbf{v}}}$ are independent Wiener processes with $\bm{\sigma}_{1}\in\mathbb{R}^{d_{\mathbf{u}_{1}}\times d_{\mathbf{u}_{1}}}$ and $\bm{\sigma}_{2}\in\mathbb{R}^{d_{\mathbf{v}}\times d_{\mathbf{v}}}$ as the noise coefficients. $\mathbf{v}=\bm{\varphi}(\mathbf{u}_{2})\in\mathbb{R}^{d_{\mathbf{v}}}$ is called latent states with a nonlinear encoder $\bm{\varphi}:\ \mathbb{R}^{d_{\mathbf{u}_{2}}}\mapsto\mathbb{R}^{d_{\mathbf{v}}}$. It is worth noting that the conditional distribution of the latent variables $p(\mathbf{v}(t)|\mathbf{u}_{1}(s),s\leq t)$ is Gaussian, due to the conditional linear structure in Eq. (2.8), which facilitates efficient analytic formulae of DA with more details discussed in Section 2.2.2. With the state prediction and DA results of $[\mathbf{u}_{1};\mathbf{v}]$, the corresponding original unobserved state variables can be obtained via $\mathbf{u}_{2}=\bm{\psi}(\mathbf{v})$ with the nonlinear decoder $\bm{\psi}:\ \mathbb{R}^{d_{\mathbf{v}}}\mapsto\mathbb{R}^{d_{\mathbf{u}_{2}}}$. The temporal derivative of the system states $\mathrm{d}\mathbf{u}/\mathrm{d}t$ is sometimes denoted as $\dot{\mathbf{u}}$ for short. The encoder $\bm{\varphi}$, decoder $\bm{\psi}$, and the four nonlinear functions $\mathbf{f}_{1}$, $\mathbf{g}_{1}$, $\mathbf{f}_{2}$, and $\mathbf{g}_{2}$ are parameterized by neural networks and are jointly learned from data $\{\mathbf{u}^{\star}(t_{n})\}_{n=0}^{N}$. The neural networks that approximate the four nonlinear functions are collectively referred to as $\bm{\eta}$.
The CGKN comprises an encoder $\bm{\varphi}$, a decoder $\bm{\psi}$, and sub-networks $\bm{\eta}$. More specifically, the architecture of CGKN consists of (i) an encoder that maps the original unobserved system state variables to the latent state variables of a conditional Gaussian nonlinear system, (ii) neural networks that approximate the unknown terms in the dynamics of the conditional Gaussian nonlinear system, and (iii) a decoder that transforms the latent state variables of the conditional Gaussian nonlinear system back to the original unobserved system state variables. The architecture of CGKN is illustrated in Fig. 2.2(a). As a high-level deep learning framework, CGKN can integrate various neural network architectures, including fully-connected neural networks, residual neural networks, convolutional neural networks, and recurrent neural networks for encoder $\bm{\varphi}$, decoder $\bm{\psi}$, and sub-nets $\bm{\eta}$. In this work, the fully-connected neural networks are used to demonstrate the effectiveness and capability of CGKN.
It is worth noting that the CGKN framework is also applicable to another general type of a partially observed system, which consists of a dynamical model for the whole system state and an observation mapping that partially extracts observable information from the system state. This type of partially observed system can be formalized in the same form as shown in Eq. (LABEL:eq:DS_u1u2) and then modeled by the CGKN in Eq. (2.8). This is achieved by retaining the dynamical model and modeling the dynamics of the observations in relation to the system states, where the system state corresponds to the unobserved states $\mathbf{u}_{2}$ and the observations correspond to the observed states $\mathbf{u}_{1}$ in Eq. (2.8). If the dynamics of observations do not exist or if the observation time interval is large, a discrete-time version of CGKN can be employed.
@@
Figure 2.2: The architecture of CGKN and its applications for state prediction and data assimilation. The CGKN approximates the true nonlinear complex dynamical system by a modeled system with a conditional linear structure, which is also known as a conditional Gaussian nonlinear system. The conditional linear structure is enabled by a proper choice of latent state variables $\mathbf{v}$ as illustrated in (a). The nonlinear mappings between the original state variables $\mathbf{u}_{2}$ and the latent state variables $\mathbf{v}$ are achieved via the encoder $\bm{\varphi}$ and decoder $\bm{\psi}$, which are jointly learned with the unknown functions $\mathbf{f}_{1}$, $\mathbf{g}_{1}$, $\mathbf{f}_{2}$, $\mathbf{g}_{2}$ of the conditional Gaussian nonlinear system. The pre-trained CGKN can be used for state prediction and efficient data assimilation as illustrated in (b) and (c).
@@ 2.2.1 CGKN for State Prediction
The CGKN model can be used for state prediction given an initial state $\mathbf{u}^{\star}(t_{0})=[\mathbf{u}_{1}^{\star}(t_{0});\mathbf{u}^{\star}_{2}(t_{0})]$. Firstly, the unobserved state variables $\mathbf{u}^{\star}_{2}(t_{0})$ are transformed into latent state variables $\mathbf{v}^{\star}(t_{0})$ via an encoder $\bm{\varphi}$, which leads to the initial state of conditional Gaussian nonlinear system $\mathbf{u}_{\textrm{CG}}=[\mathbf{u}_{1};\mathbf{v}]$ for Eq. (2.8):
$\mathbf{u}_{\textrm{CG}}^{\star}(t_{0})=\begin{bmatrix}\mathbf{u}^{\star}_{1}(t_{0})\\ \mathbf{v}^{\star}(t_{0})\end{bmatrix}=\begin{bmatrix}\mathbf{u}^{\star}_{1}(t_{0})\\ \bm{\varphi}(\mathbf{u}^{\star}_{2}(t_{0}))\end{bmatrix}.$ (2.9)
@@
$\mathbf{u}_{\textrm{CG}}(t_{n})=\begin{bmatrix}\mathbf{u}_{1}(t_{n})\\ \mathbf{v}(t_{n})\end{bmatrix}=\begin{bmatrix}\mathbf{u}^{\star}_{1}(t_{0})+\int_{t_{0}}^{t_{n}}(\mathbf{f}_{1}(\mathbf{u}_{1}(t))+\mathbf{g}_{1}(\mathbf{u}_{1}(t))\mathbf{v})\mathrm{d}t\\ \mathbf{v}^{\star}(t_{0})+\int_{t_{0}}^{t_{n}}(\mathbf{f}_{2}(\mathbf{u}_{1}(t))+\mathbf{g}_{2}(\mathbf{u}_{1}(t))\mathbf{v})\mathrm{d}t\end{bmatrix}.$ (2.10)
@@
In this work, the encoder $\bm{\varphi}$, the decoder $\bm{\psi}$, and the functions $\mathbf{f}_{1}$, $\mathbf{g}_{1}$, $\mathbf{f}_{2}$, $\mathbf{g}_{2}$ are jointly learned to ensure a good state prediction performance, i.e., the difference between the state prediction $\mathbf{u}(t_{n})$ and the true system state $\mathbf{u}^{\star}(t_{n})$ is small for a series of $t_{n}$. This is achieved by incorporating the state prediction performance into the loss function, and more details of the loss function for training a CGKN model are summarized in Section 2.3. An illustration of CGKN for state prediction is shown in Fig. 2.2(b).
If the initial value of unobserved states $\mathbf{u}^{\star}_{2}(t_{0})$ are not given, the estimated unobserved states via data assimilation after a warm-up period can substitute for the unknown true initial value. The CGKN for DA is introduced in Section 2.2.2. Forecasting the entire system state using data from partially observed states is one of the features of CGKN, and it is worth noting that the performance of the state forecast will thus depend on the quality of DA estimated unobserved states.